OpenStudy (anonymous):

What is the oblique (slant) asymptote of the function f(x) = the quantity of x squared plus 7 x plus 11, all over x plus 5? Help me please.

OpenStudy (anonymous):

\[\frac{ x^2+7x+11 }{ x+5 }\]

OpenStudy (anonymous):

Answer Choices: A. y = x + 1 B. y = x − 2 C. y = x + 4 D. y = x + 2

OpenStudy (anonymous):

@tom982 @phi help me?

OpenStudy (anonymous):

@Hero @dan815 @Jaynator495

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (phi):

divide x+5 into the top expression

OpenStudy (phi):

you will get \[f(x)= x+2 + \frac{1}{x+5} \] (the last term comes from writing the remainder over the divisor) for x very big, 1/(x+5) becomes a small number, and you have (approximately) the equation y=x+2 (though it never exactly equals that because 1/(x+5) only gets close to zero, it never becomes zero)

OpenStudy (anonymous):

Thank you.